Find the wavelength associated with Virar-f local train having mass 100x10^3 kg moving with the speed of 23.76 km/hr. ( plank’s constant= 6.6x 10^-34 Js)

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
According to De Broglie proinciple 
wavelength = h/P , where  P = m × V
 M=  mass,
V = velocity = 23.76 km/hr
V = 23.76× 1000/3600 m/s
V= 6.6 m/s
wavelength =(6.6 × 10-34)/ (100 ×103)×6.6
 wavelength = 1 ×10-39 m

wavelength =  10-29 

Regards
Arun (askIITians forum expert)